a toy manufacturer uses this acceptance sampling plan: Randomly select and test 47 batteries and determine what batteries do not meet specifications. A shipment contains 4000 batteries, and 1% of them do not meet specifications. What is the prob be accepted , or will many be rejected?


1
Expert's answer
2021-03-25T02:25:59-0400

Solution:

Let "X" be the number of defective batteries.

The probability distribution of the random variable "X", hypergeometric distribution, is

"h(x ; n, M, N)=\frac{\left(\begin{array}{l}
M \\
x
\end{array}\right)\left(\begin{array}{l}
N-M \\
n-x
\end{array}\right)}{\left(\begin{array}{l}
N \\
n
\end{array}\right)}"

Given: "N=4000, M=4000 \times 0.01=40, n=47" .

This part is missing, so we assume we are asked "P(X\le 3)".

"P(X\le3)=P(X=0)+P(X=1)+P(X=2)+P(X=3)"

"\begin{array}{c}
=\frac{\left(\begin{array}{c}
40 \\
0
\end{array}\right)\left(\begin{array}{c}
4000-40 \\
47-0
\end{array}\right)}{\left(\begin{array}{c}
4000 \\
47
\end{array}\right)}+\frac{\left(\begin{array}{c}
40 \\
1
\end{array}\right)\left(\begin{array}{c}
4000-40 \\
47-1
\end{array}\right)}{\left(\begin{array}{c}
4000 \\
47
\end{array}\right)}+ \\
+\frac{\left(\begin{array}{c}
40 \\
2
\end{array}\right)\left(\begin{array}{c}
4000-40 \\
47-2
\end{array}\right)}{\left(\begin{array}{c}
4000 \\
47
\end{array}\right)}+\frac{\left(\begin{array}{c}
40 \\
3
\end{array}\right)\left(\begin{array}{c}
4000-40 \\
47-3
\end{array}\right)}{\left(\begin{array}{c}
4000 \\
47
\end{array}\right)} 
\end{array}"

"\approx 0.62181+0.29867+0.06843+0.00996 \\
\approx 0.99887"